OpenStudy (anonymous):

Solve, using the substitution method. y = 3x + 5 4x – y = 5

OpenStudy (anonymous):

A. (10, 35) B. (15, 10) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

A opt is the answer, what grade are you studying in?

OpenStudy (anonymous):

y= 3x+5.....(1) 4x - y=5.....(2) put the value of y in 2 we get: 4x - ( 3x+5) = 5 => 4x-3x-5=5 =>x=5+5=10 now put x=10 in (1) or (2) y=3(10)+5=35 thus point is (10,35)